52 68 Problem 6. (8 points) A six-sided die is loaded in a way that each odd number is twice as likely as each even number. All odd faces are equally likely, as are all even faces. (1) (5 points) Construct a probabilistic model for a single roll of this die; By constructing a probablistic model it means you should provide the sample space and the probabilities of all single events. (2) (3 points) Find the probability that the outcome is less than 4.
Answer #1

Solution:-

6)

i)

P(x) 0.22222 0.22222 0.22222 2 4 5 6 Sum

ii) P(x < 4) = 0.5555

P(x < 4) = P(x = 1) + P(x = 2) + P(x = 3)

P(x < 4) = 0.2222 + 0.11111 + 0.2222

P(x < 4) = 0.5555